OpenStudy (anonymous):

The height of a parallelogram is 2 units more than the base. The measure of the base of the parallelogram is 10 units. What is the area, in square units, of the parallelogram? Answer 60 100 120 200

OpenStudy (anonymous):

a=b*h=(2+10)(10)=120

OpenStudy (anonymous):

Good job!!

OpenStudy (anonymous):

thanks